There are three forces acting on a uniform beam of 5 kg.F1=200N,F2=80N and the system is in equilibrium.Find x if l=4cm and give the magnitude of F3?
1
Expert's answer
2018-05-07T09:24:08-0400
F_1=F_2+f_3
F_3=200-80=120 N.
Taking moments:
200(4-x)=120x
x=2.5 cm.